OpenStudy (anonymous):

HELP ASAP! The us senate has 100 members, all democrats or Republicans. Recently there were 12 more Democrats than Republicans. How many senators from each political party were there at that time? STEPS NEEDED! :)

OpenStudy (anonymous):

The u.s. senate has 100 members, all democrats or repulbicans. Recently there were 12 more Democrats than Republicans. How many senators from each political party were there at that time? Choices for the number of republicans:38, 42, 44. ---------------------------- Quantity Equation: D + R = 100 Quantity Equation: D = R+12 ----------------------------------- Substitute to solve for "R": R+12 + R = 100 2R = 88 R = 44 (# of Republicans) ============================

OpenStudy (anonymous):

hopes this helps

OpenStudy (anonymous):

THANKS!